# Repeating nuclear transients from repeating partial tidal disruption events: reproducing ASASSN-14ko and AT2020vdq

### v2

@@ 4.3 Evolution of the Orbital Period and the Peak Fallback Timescale
Our analytical estimates from Section 2 show that for solar-like stars being repeatedly tidally disrupted by an SMBH, the angular velocity imparted to the stellar core can change the period by a small fraction of its original value. For the specific case of ASASSN-14ko, having a period of $114\pm 1$ days, we estimated the fractional change in the period to be $\Delta T/T\simeq-0.003$, which is in excellent agreement with observations (Payne et al., 2023). We note, however, that to continue to reduce the orbit by the same fractional amount, this energy must be lost from the star and efficiently, i.e., on a per-orbit basis. Since the imparted angular momentum to the star results in an velocity that is approximately uniform, it is unlikely that this energy will be dissipated efficiently, meaning that – if the imparted rotation does represent the sink of orbital energy that leads to the period decay – one would expect a pronounced $\ddot{P}$ as the star is not efficiently torqued above $\Omega_{\rm p}$.
In addition to bulk rotation, one expects a comparable amount of energy produced in oscillatory modes that is ultimately dissipated as heat, either viscously or through three-mode (and higher) nonlinear couplings (e.g., McMillan et al. 1987; Kochanek 1992; Kumar & Goodman 1996; Weinberg et al. 2012). As we discuss in Appendix A, we expect most of the thermal energy to be concentrated in the outer layers of the star where the amplitude of the tidal acceleration is largest, and these layers are removed on subsequent encounters with the SMBH. For ASASSN-14ko (i.e., with an orbital period of 114 days and a SMBH mass of $10^{7}M_{\odot}$), the reduction in the orbital period would result in the deposition of $\sim 1.5\times 10^{46}$ erg in $0.01M_{\odot}$ (i.e., comparable to the amount of mass stripped from the star in our simulations), which would require a luminosity of $L\approx 3.5L_{\rm Edd}$ if this energy were to be lost radiatively ($L_{\rm Edd}$ is the Eddington luminosity of a $3M_{\odot}$ star, assuming $\kappa_{\rm es}=0.34$ cm2 g-1 for the opacity; see the discussion in Appendix A for more details). The amount of energy dissipated via tides and required to reproduce the observed $\dot{P}$ therefore could not be efficiently exhausted radiatively (see also Kumar & Quataert 1996), but the tidal stripping of the outer layers of the star serves an alternative outlet through which the energy can be lost mechanically and without destroying the high-density (and thermally insulated; again, see Appendix A) core, thus allowing periodic accretion events to continue. While our resolution tests and the implementation of alternative thermodynamic treatments, discussed in detail in Appendix A, provide evidence to suggest that this is the means by which the star can continue to survive many tidal encounters with the SMBH, additional investigations – particularly in the context of linear tidal theory – are required to solidify this connection.
@@ 5 Summary
    Using the energy-period relation for the Keplerian orbit, we showed that the rotational velocity imparted to the surviving stellar core can lead to changes in its period that are comparable to the observed period derivative of $\dot{P}=-0.0026$ for ASASSN-14ko. Once the star is spun up to $\sim\Omega_{\rm p}$, if additional energy is imparted into oscillatory and quadrupolar oscillations that can be subsequently lost, this effect would continue on a per-orbit basis. Our arguments (Section 4.3) and simulations (Appendix A) suggest that the mechanism responsible for removing the energy is not predominantly radiative. Instead, it is likely (and consistent with the results in Appendix A) that the thermal energy is deposited in the outer layers of the star that are tidally stripped on subsequent encounters, while the core is relatively unmodified and allowed to survive for mamy more encounters.

### v3

@@ 4.3 Evolution of the Orbital Period and the Peak Fallback Timescale
Our analytical estimates from Section 2 show that for solar-like stars being repeatedly tidally disrupted by an SMBH, the angular velocity imparted to the stellar core can change the period by a small fraction of its original value. For the specific case of ASASSN-14ko, having a period of $114\pm 1$ days, we estimated the fractional change in the period to be $\Delta T/T\simeq-0.003$, which is in excellent agreement with observations (Payne et al., 2023). We note, however, that to continue to reduce the orbit by the same fractional amount, this energy must be lost from the star and efficiently, i.e., on a per-orbit basis. Since the imparted angular momentum to the star results in an angular velocity that is approximately uniform, it is unlikely that this energy will be dissipated efficiently, meaning that – if the imparted rotation does represent the sink of orbital energy that leads to the period decay – one would expect a pronounced $\ddot{P}$ as the star is not efficiently torqued above $\Omega_{\rm p}$.
In addition to bulk rotation, one expects a comparable amount of energy produced in oscillatory modes that is ultimately dissipated as heat, either viscously or through three-mode (and higher) nonlinear couplings (e.g., McMillan et al. 1987; Kochanek 1992; Kumar & Goodman 1996; Weinberg et al. 2012). As we discuss in Appendix A, we expect most of the thermal energy to be concentrated in the outer layers of the star where the amplitude of the tidal acceleration is largest, and these layers are removed on subsequent encounters with the SMBH. For ASASSN-14ko (i.e., with an orbital period of 114 days and a SMBH mass of $10^{7}M_{\odot}$), the reduction in the orbital period would result in the deposition of $\sim 1.5\times 10^{46}$ erg in $0.01M_{\odot}$ (i.e., comparable to the amount of mass stripped from the star in our simulations), which would require a luminosity of $L\approx 3.5L_{\rm Edd}$ if this energy were to be lost radiatively ($L_{\rm Edd}$ is the Eddington luminosity of a $3M_{\odot}$ star, assuming $\kappa_{\rm es}=0.34$ cm2 g-1 for the opacity; see the discussion in Appendix A for more details). The amount of energy dissipated via tides and required to reproduce the observed $\dot{P}$ therefore could not be efficiently exhausted radiatively (see also Kumar & Quataert 1998), but the tidal stripping of the outer layers of the star serves an alternative outlet through which the energy can be lost mechanically and without destroying the high-density (and thermally insulated; again, see Appendix A) core, thus allowing periodic accretion events to continue. While our resolution tests and the implementation of alternative thermodynamic treatments, discussed in detail in Appendix A, provide evidence to suggest that this is the means by which the star can continue to survive many tidal encounters with the SMBH, additional investigations – particularly in the context of linear tidal theory – are required to solidify this connection.
@@ 5 Summary
    Using the energy-period relation for the Keplerian orbit, we showed that the rotational velocity imparted to the surviving stellar core can lead to changes in its period that are comparable to the observed period derivative of $\dot{P}=-0.0016$ for ASASSN-14ko, provided the star is spun up to a rotational velocity comparable to $\sim\Omega_{\rm p}$ in one orbit. Our simulations show (see Figure 5), however, that it takes several pericenter passages to spin up the star by this amount, and thus the rotational energy of the core does not act as an efficient sink for the dissipated orbital energy (at least as concerns the observed $\dot{P}$ for ASASSN-14ko). Moreover, the per-orbit increment in $\Omega$ declines with time and the star can not be spun up beyond $\sim\Omega_{\rm p}$, implying that there would be a significant $\ddot{P}$ if this were the primary source of dissipation. Instead, it is likely (and consistent with the results in Appendix A) that tidal dissipation deposits thermal energy (at the expense of the binding energy of the orbit) in the outer layers of the star that are tidally stripped on subsequent encounters, i.e., the energy is lost mechanically instead of radiatively, while the core is relatively unmodified and survives many more encounters.
@@ Appendix A Effects of tidal heating
Finally, we note that the degree to which tidal dissipation modifies the orbital period of the star is a nontrivial and non-monotonic function of the energy. In particular, while tides dissipate orbital energy and further bind the star to the black hole for sufficiently small $\beta$, higher-order moments in the black hole’s gravitational field can impart a positive energy to the surviving core once the star loses a sufficiently large amount of mass, as shown by Manukian et al. (2013); Gafton et al. (2015); Cufari et al. (2023) (see Kremer et al. 2022 for an analysis of this phenomenon for stellar-mass black hole encounters). However, these investigations were specific to the case where the star was originally on a parabolic orbit, was not rotating, and had a simple (i.e., polytropic) density profile, none of which apply to the tidal encounters considered here. While we find – by following the same procedure as outlined in Cufari et al. (2023) – that the $3M_{\odot}$ TAMS star on a $\beta=1$ orbit experiences a reduction in its orbital energy at the level of $0.01GM_{\star}/R_{\star}$, such that its orbital period would be reduced by an amount that is in rough agreement with the $\dot{P}$ exhibited by ASASSN-14ko (see the preceding paragraph, but we emphasize that our orbit was parabolic, and it is not clear how an initially bound orbit would modify this result), we defer a detailed investigation of this aspect of this problem to future work.